Consider two streams of cash flows, A and B. Stream A’s first cash flow is $9,500 and is received three years from today. Future cash flows in Stream A grow by 4 percent in perpetuity. Stream B’s first cash flow is −$9,400, is received two years from today, and will continue in perpetuity. Assume that the appropriate discount rate is 12 percent.

a.

What is the present value of each stream? (A negative amount should be indicated by a minus sign. Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

b.

Suppose that the two streams are combined into one project, called C. What is the IRR of Project C? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

1.
A=9500/1.12^3*1/(1-1.04/1.12)=94666.77296

B=-9400/(1.12*0.12)=-69940.47619

2.
9500/(1+IRR)^3*1/(1-1.04/(1+IRR))-9400/((1+IRR)*IRR)=0
=>IRR=22.6915%